IDENTIFICATION OF NONLINEARITY IN CONDUCTIVITY EQUATION VIA DIRICHLET-TO-NEUMANN MAP

HYEONBAE KANG GEN NAKAMURA

Abstract. We prove that the linear term and quadratic nonlinear term entering a nonlinear elliptic equation of divergence type can be uniquely identified by the Dirichlet to Neuman map. The unique identifiability is proved using the complex geometrical optics solutions and singular solutions.